OpenStudy (anonymous):

integration for : 1/(1+e^x)

OpenStudy (mr.math):

Hint: \[\frac{1}{1+e^x}=\frac{1+e^x-e^x}{1+e^x}=1-\frac{e^x}{1+e^x}.\]

OpenStudy (anonymous):

ok then what will i have to do after that ??

OpenStudy (mr.math):

I'm sure you know how to integrate 1. For the term \(\frac{e^x}{1+e^x}\), use a substitution \(u=1+e^x\). It's an easy one.

OpenStudy (anonymous):

sorry mr.math tell me slowly the steps to final im studying from 19 hrs i think my brain stopped

OpenStudy (anonymous):

aha i integrated it

OpenStudy (mr.math):

Alright! So we have so far: \[\int \frac{1}{1+e^x}dx=\int (1-\frac{e^x}{1+e^x})dx=\int dx-\int \frac{e^x}{1+e^x}dx.\]

OpenStudy (anonymous):

thank you very much it will be ( x - in ( 1+e^x ) + c right ?

OpenStudy (mr.math):

Exactly! :)

OpenStudy (anonymous):

thank you for caring

OpenStudy (mr.math):

You're welcome.